Attendees: J. Acree, R. Fisher, R. McCullough, C. Niday, D. Wirt
Denise reported on the completion of the office moves. Carla, Wes and Denise are now in their new offices. Denise has assumed responsibility for performing all propagation studies and Wes is working with the more technical aspects of the NRQZ. Denise expressed her appreciation for everyone’s input regarding web site ideas and she will be working to update and improve the web site as the NRQZ workload permits.
This item will be reported on during next month’s IPG meeting.
Denise reported that the GBT control room shielding rework has been completed. This includes the four 48” x 48” windows, the patio door with two large windows and the interior door leading to the hallway. Wes and Nathan are working on the other interior doors adjacent to the hallway. Jeff indicated that work on the OVLBI control room windows is on hold until a later date. Out of consideration for the observers, Jeff is reluctant to pull the temporary protection from the large windows should an unforeseen delay occur as it did while working on the windows in the GBT control room. In an effort to solve the cold weather condensation problem, Butch Wirt designed a prototype cover to provide a thermal break on the shielded room windows. The covers will be installed on the interior side of the windows and secured to the windows using standoffs. This will allow easy removal and will not cause any penetrations to the shielding. Fans in the GBT equipment room and plastic covering in the GBT control room are in place as temporary measures. The humidity is now stabilized at approximately 45-50%. It was noted that none of the window contractors that Butch spoke with were willing to guarantee an exterior storm window to control condensation.
Jeff reported several members of the IPG and receiver group have been working on a RFI problem associated with the LO receiver module in the GBT receiver room LO rack. The problem has been loosely described as a “5 MHz comb” from 500 MHz to several GHz. The strength of the comb was reduced by about 30 dB by disabling some fault monitoring circuitry and adding some shielding, but Mike Stennes feels the LO receiver module and its spare should be rebuilt when time permits.
Jeff reported that Spectrum Control, Inc. resin sealed PI filters, such as those that were used in the AEVC HVAC system, must be installed per manufacturers recommendations. For example, they should be tightened using only 4-6 in-lbs of torque and should be soldered using a temperature controlled soldering iron set to 525 " 10 EF, and even then, a heat sink should be attached next to the filter body during soldering. In the AEVC installation, he believes none of this occurred. After a period of time, major failures such internal shorting and self destruction were experienced in a good number of filters. It appears the failures were triggered by voltage spikes that were well within the manufacturing recommended operating range, and per the manufacturer, such failures are symptomatic of improper installation. Some of the failed units were sent back to the manufacturer for analysis, but the findings were inconclusive. Jeff believes improper installation to be the culprit and will organize a replacement project after the AEVC is turned over to NRAO. Hundreds of filters are involved.
Based on tests he performed with the laser range finders, Randy suggested considering the use of conductive epoxy during the replacement installation. Similarly, Chuck stated another possibility would be to thread the filter through a piece of copper tape during installation. Randy will prepare an installation document for the filter replacement project.
Randy gave an update on RFI testing and control measures in support of PTCS. No chamber testing had been conducted, but Randy indicated that two boxes were nearly ready for testing. One box will have his full RFI measures in place and the other box without any RFI control measures. This is being done at the request of Kim Constatikes to establish the value of the RFI control.
Randy will bring the “homemade” continuum receiver back into use. In doing so, it will need an RF filter to be able to select the bands, as well as a remote switch for on/off testing. The computer and continuum receiver are reported to work fine, although the computer is a 486 machine. Tim Weadon and Rich Lacasse are the designers, and would be the resources for information. The intent of this set up was to improve sensitivity and reduce test time. To everyone’s knowledge, it works just fine, but is better suited for use with the right application. Randy expects to report actual test results by next meeting.
Rick inquired about what information we have available to identify noise sources by frequency. His question was prompted by work being performed by Karen O’Neil in the 1150 – 1250 MHz band. Databases we have include the FCC master frequency database (mid-Atlantic region) as well as the government database. One area in which we lack information is satellites. Rick believes that the commissioning task devoted to this will fall short of supplying readily available information needed by the Observers. Jeff agreed this is an important issue and will present a schedule of projects next IPG meeting. Jeff will also have Wes summarize the steps he takes in RFI source identification.